Hello! This was a bit confusing questions for me. I was thinking of changing the percents to decimals. But would I add them together for all friends that show up to the party? Thank you for your help! I am really stuck on this one
Suppose you are planning to throw a party this weekend and invite 5 friends. Knowing your friends’ history, some of them tend to cancel at the last minute. Mike has shown up to 95% of the time, Jamie has shown up 80% of the time, Jack shows up 75% of the time, Aimee shows up 85% of the time, and Kris shows up 50% of the time. Suppose that a chance an individual friend shows up is independent.
Find the probability that all the friends show up to the party.

Find the probability that no of the friends up to the party.

Answer by ikleyn(52775)

.
Suppose you are planning to throw a party this weekend and invite 5 friends. Knowing your friends’ history,
some of them tend to cancel at the last minute. Mike has shown up to 95% of the time, Jamie has shown up 80% of the time,
Jack shows up 75% of the time, Aimee shows up 85% of the time, and Kris shows up 50% of the time.
Suppose that a chance an individual friend shows up is independent.


(a)   Find the probability that all the friends show up to the party.

         P = 0.95*0.80*0.75*0.85*0.50 = use your calculator.


         You simply need to multiply all relevant probabilities 

             (because the events are independent).


(b)   Find the probability that no of the friends up to the party.

        P = (1-0.95)*(1-0.8)*(1-0.75)*(1-0.85)*(1-0.50) = use your calculator.



         Again, you simply need to multiply all relevant probabilities 

             (because the events are independent).

         This time the relevant probabilities are the COMPLEMENTS (!) to the given probabilities.

Solved.

-------------

To better learn the subject,  read this lesson
    - Solving probability problems using complementary probability
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Solved problems on Probability".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.